OpenStudy (anonymous):

A coil with large distribution capacitance has a: (i) high Q , (ii) low Q , (iii) low resonant frequency , (iv) high resonant frequency.

OpenStudy (festinger):

Resonant frequency is: \[\omega=\frac{1}{\sqrt{LC}}\] Since\[\omega=2\pi f\] \[f=\frac{1}{2\pi\sqrt{LC}}\] So a larger C gives a smaller resonant frequency. Charge Q is given by the relation Q=CV. For the same V, a larger C will give a larger Q. Thus (i) and (iii)